OpenStudy (anonymous):

I need help in geometry. @jim_thompson5910 @HawkCrimson @Mertsj

jimthompson5910 (jim_thompson5910):

ok what's the question

OpenStudy (anonymous):

1 second let me get it all in pictures because its kinda big..

OpenStudy (anonymous):

i love that i was classified with @jim_thompson5910 and @Mertsj when im not even close to their level :D

OpenStudy (anonymous):

ok this is the first part: http://oi50.tinypic.com/33wb05z.jpg this is the second: http://oi47.tinypic.com/v6k9xi.jpg

jimthompson5910 (jim_thompson5910):

you don't give yourself enough credit

OpenStudy (anonymous):

thx :D

OpenStudy (anonymous):

i asked someone for help and they sent me your name guys, i hope you dont mind

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (anonymous):

i have no idea how to do it? like for the questions that i have to make up

jimthompson5910 (jim_thompson5910):

well if you know 2 sides of a right triangle, how can you find the third?

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

you add them and substract that from 180

jimthompson5910 (jim_thompson5910):

you're probably thinking angles

jimthompson5910 (jim_thompson5910):

but no you would use the pythagorean theorem

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2

jimthompson5910 (jim_thompson5910):

plug in the known values and solve for the third unknown value

OpenStudy (anonymous):

A^2 + B^2 = C^2

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so thats the answer for that part?

jimthompson5910 (jim_thompson5910):

well you could go into a bit more detail if you wanted to, but that's the basic answer

OpenStudy (anonymous):

am good with that, ill use that, how about for the second one now?

jimthompson5910 (jim_thompson5910):

for #2, the answer is yes, but only if you know basic trig (like sine, cosine, etc)

jimthompson5910 (jim_thompson5910):

if you know how all that works, then you can if not, then you won't have the proper tools

OpenStudy (anonymous):

what would i have to use for finding the hypothenuse?

OpenStudy (anonymous):

SohCahToa. so u can use sin or Cos

OpenStudy (anonymous):

so how do i explain the answer for the second one in the first picture?

OpenStudy (anonymous):

how do i put it into words?

jimthompson5910 (jim_thompson5910):

well for instance, if you know one angle and the opposite side, then you can use the sine function to find the hypotenuse since sine = opposite/hypotenuse

OpenStudy (anonymous):



OpenStudy (anonymous):

u can use that as an example

OpenStudy (anonymous):

ok so it would be sine = 5/39

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh thats for the hypothenuse my bad

OpenStudy (anonymous):

sin39 = 5/h

OpenStudy (anonymous):

yep :) cool so u do know it

OpenStudy (anonymous):

a little =/

OpenStudy (anonymous):

 how do u solve for x?

OpenStudy (anonymous):

is it the same way you put it before? like changing the numbers, would it be sin 42 = 15h?

OpenStudy (anonymous):

15/h

OpenStudy (anonymous):

think about it, sin( taeta ) = opp/ hyp wat is the side opp the angle and wat is the hypotenus?

OpenStudy (anonymous):

am still lost, this is the last assigment in this section, can u tell me how to do the two questions remaining please? :( its 3 nd 4

OpenStudy (anonymous):

sry os is being stupid

OpenStudy (anonymous):

ill give u a hint since u seem smart enough :) Sin isnt ur only choice, u can also use Cos and Tan

OpenStudy (anonymous):

and ok ill just help with 3 and 4 then, moving on

OpenStudy (anonymous):

u there @intronic ?

OpenStudy (anonymous):

yes sorry am here